OpenStudy (anonymous):

y^3+13^2+40y is this prime

OpenStudy (raden):

because cant be factored, so yes it is A prime polynomial

OpenStudy (anonymous):

THAK YOU AGAINF CLARITY

OpenStudy (raden):

you're welcome...

OpenStudy (anonymous):

x^2-x-30 I have no idea how to factor this because there is only 1 whole number so does that make it a prime also?

OpenStudy (raden):

no, its not a prime... because it can be factored x^2-x-30 = (x-6)(x+5)

OpenStudy (anonymous):

okay so there does not have to be two numbers

OpenStudy (raden):

ohhh... u have remembered my coment before ^^ i meant, if u find 2 numbers when u added be -1 and when multiplied be -30, that numbers are -6 and 5, right ? so. the factor form of it = (x-6)(x+5)

OpenStudy (anonymous):

wow I wish this came as easy to me